There are two more quarters than dimes, and as many nickels, as quarters and dimes together. The total amount of money is $3.75. How many quarters, dimes, and nickels are there?
Answer by Boreal(15235)

dimes=x
quarters=x+2
nickels=2x+2
sum is 3.75
.10x+.25(x+2)+0.05(2x+2)=3.75
.10x+.25x+.50+0.10x+0.10=3.75
0.45x=3.15
x=7 dimes (0.70)
x+2=9 quarters (2.25)
2x+2=16 nickels (0.80)
They add to $3.75